### v3

# TinkerXR: In-Situ, Reality-Aware CAD and 3D Printing Interface for Novices

Abstract.
Despite the growing accessibility of augmented reality (AR) for visualization, existing computer-aided design (CAD) systems remain confined to traditional screens or require complex setups or predefined parameters, limiting immersion and accessibility for novices. We present TinkerXR, an open-sourced interface enabling in-situ design and fabrication through Constructive Solid Geometry (CSG) modeling. TinkerXR operates solely with a headset and 3D printer, allowing users to design directly in and for their physical environments. By leveraging spatial awareness, depth occlusion, recognition of physical constraints, reference objects, and intuitive hand movement controls, TinkerXR enhances realism, precision, and ease of use. Its AR-based workflow integrates design and 3D printing with a drag-and-drop interface for a 3D printer’s virtual twin. A user study comparing TinkerXR with TinkerCAD demonstrates higher accessibility, engagement, and ease of use for novices. By bridging the gap between digital creation and physical output, TinkerXR transforms everyday spaces into accessible and expressive creative studios.
computer-aided design, computer-aided manufacturing, mixed reality, augmented reality, extended reality, 3D printing, situated design
journalyear: 2025copyright: rightsretainedconference: arXiv; preprint; 2025booktitle: arXiv
Figure 1. TinkerXR is an in-situ CAD and 3D printing AR interface. The figure illustrates key features such as spatial awareness, depth occlusion, and the use of physical objects and constraints for design. With Constructive Solid Geometry modeling, users can convert surfaces into interactive design canvases, and after designing, interact with 3D printers’ digital twins to prepare and initiate fabrication. The example designs include a pen holder, phone stand, cable organizer, plant pot holder, bottle mount, and key hook.
[…]
Five images labeled (a), (b), (c), (d), and (e) show the user interacting with the system. In pictures (a), (b), and (c), the user is seen wearing a headset. In picture a, the user is in a room with a desk. The user is standing up next to the desk and pointing their finger at the desk. A ray from their finger hits the table, where a text appears on top, saying ”select: table”. Three similar lines are extended towards the floor, wall, and ceiling. In pictures (b) and (c), the user is sitting on the desk, where a grid is seen covering the top surface of the desk. In the picture (b), the user is creating a cube object with a poking gesture with their right hand on the menu on the inside of their left wrist. In picture (c), the user is holding a real-world finger skateboard on their left hand while applying a pinch gesture with a right hand onto a virtual finger skateboard obstacle. In picture (d), the first-person view of the user is seen where a virtual printer twin is on top of the desk and the user is dragging the designed model with their left hand on the printer twin. The user is poking the button next to the digital printer twin labeled ”Start print”. In picture (e), the 3D-printed result can be seen on top of the desk.
1. Introduction
Augmented reality (AR) has emerged as a transformative technology across various fields, notably in design and manufacturing. Traditional computer-aided design (CAD) tools such as SolidWorks and Blender have provided high levels of precision and functionality (Reyes, 2018; Mullen, 2012), while platforms like Tinkercad have made CAD more accessible to novices and educational users (Jacob et al., 2021; Mahapatra et al., 2019). Despite the advances in AR, however, these tools are still mostly limited to traditional 2D screens, which limits their interactivity and integration with the user’s real-world environment.
The introduction of Tinkercad AR [^1] https://www.tinkercad.com/ipad-app, Accessed: 16.07.2024 marked a notable progression by incorporating AR to view designs in physical spaces, albeit mainly for visualization. Applications such as Gravity Sketch [^2] https://www.gravitysketch.com/, Accessed: 16.07.2024 have begun to leverage virtual reality (VR) headsets and controllers to enable more immersive 3D sketching experiences. In recent HCI research, several notable advancements have emerged to enhance spatial and intuitive design interfaces. pARam (Stemasov et al., 2024) suggests that in-situ parametric design and customization can be effectively achieved through hand interactions and environmental considerations in AR. Additionally, BrickStARt (Stemasov et al., 2023a) introduces a novel approach by integrating tangible construction blocks with mixed reality (MR) to enable users to design and test physical artifacts directly in their intended locations. ProtoAR (Nebeling et al., 2018) addresses the challenges of AR content creation by combining physical prototyping with mobile authoring tools for rapid development and interaction. Furthermore, MixFab (Weichel et al., 2014) illustrates how MR can streamline personal fabrication by facilitating direct interaction with virtual objects within an AR environment. Collectively, these advancements highlight a trend towards more spatially aware and intuitive design interfaces, which are particularly beneficial for complex tasks that integrate digital and physical spaces.
Despite these advancements, a significant gap remains in harnessing AR for interactive design that directly integrates with the user’s physical environment, balances expressivity and intuitiveness, supports seamless transitions from 3D modeling to 3D printing, and delivers all these capabilities in one cohesive package. Focusing on this gap, we introduce TinkerXR, an innovative AR interface that enables in-situ CAD and 3D printing for novices. Our system bridges the gap between simple, novice-friendly tools and immersive, spatial capabilities of advanced AR technologies, offering advanced spatial interaction techniques for designing 3D models directly in the physical environment using only hand controls. TinkerXR employs intuitive, Constructive Solid Geometry modeling with primitives similar to those in Tinkercad, but enhances the design experience by integrating these elements within the user’s physical environment through AR, making the design process more intuitive and adaptable to physical constraints and possibilities, while seamlessly integrating a 3D printing workflow.
Our system utilizes a state-of-the-art headset (Meta Quest 3333https://www.meta.com/quest/quest-3/, Accessed: 16.07.2024) and its hand-gesture controls as well as passthrough features, enabling users to interact with and create 3D models directly atop their actual desks or on other flat surfaces like walls or the ground (Figure 3a). This setup not only facilitates intuitive interaction with the digital content but also allows users to accurately gauge the size and scale of their designs in relation to their immediate surroundings, enabling them to design in a way that considers and adapts to the space they’re working within.
Unlike other tools, TinkerXR is designed to recognize and adapt to real-world scene elements like furniture, enhancing its utility and applicability in everyday spaces, letting users turn any surface into their CAD workspace. Additionally, by designing around these setups, users can leverage features like snapping to grid and referencing smaller real-world objects, allowing users to design around them, making it easier to align and create models with precision, further enhancing the design process within their physical environment. Such features are critical in creating a truly immersive and intuitive design experience that seamlessly blends digital with physical.
It also incorporates depth occlusion, a technique that ensures digital objects are visually blocked by real-world objects based on their physical location and perspective, creating a more realistic and immersive experience. This is a notable advancement over other AR or VR systems, which may not achieve the same level of integration between the digital and physical worlds.
Furthermore, TinkerXR extends the concept of AR in CAD beyond digital output to include actual manufacturing processes. It integrates seamlessly with 3D printing, incorporating a drag-and-drop metaphor that allows users to intuitively prepare models for printing (Figure 3c-e). Acting as a built-in slicing interface in AR, TinkerXR prepares models for printing while providing real-time feedback through a virtual 3D printer twin, allowing users to make adjustments and visualize the output. Users can also initiate the 3D printing operation directly from the system, ensuring a seamless transition from digital design to physical creation.
To validate the effectiveness of TinkerXR, we conducted a user study (n=10) with three design tasks. In addition to TinkerXR, we compared the results with Tinkercad as ground truth, using the same three tasks to assess performance across both systems. The study also included a survey to evaluate usability, user engagement, and the learning curve. Our findings indicate that TinkerXR enhances user engagement and supports novice users in design tasks. The results suggest high usability, with users reporting ease of use and a low learning curve.
Overall, we present a novel design and fabrication system that integrates spatial and in-situ features with intuitive hand control schemes and a straightforward design logic tailored for novice users. By bridging the gap between the digital and physical worlds, our system enables seamless design-to-manufacturing workflows. Our contributions can be summarized as follows:
  • We present TinkerXR, an innovative in-situ CAD system that enables hand-based CSG modeling directly in any physical environment, tailored for novices. The system’s novel contributions include interaction methods specifically designed for CSG design in real-world settings, such as snap-to-grid, designing on top of reference objects, and intuitive controls for applying traditional CAD functionalities with hand interactions.
  • We integrate 3D printing workflows directly into the AR design process, allowing users to seamlessly transition from digital design to physical creation. The system includes a built-in slicer, real-time feedback via a virtual 3D printer twin, and the ability to initiate printing directly from the interface, offering a cohesive design-to-manufacturing pipeline.
  • We conduct a user study evaluating the effectiveness of TinkerXR when compared to Tinkercad, demonstrating high user engagement and immersiveness for novice users with no to little prior CAD experience.
  • We make TinkerXR open-source [^4] The link was removed for anonymity., contributing to the research community by providing the system and codebase for further development, experimentation, and possible integration with additional fabrication technologies.
Our approach simplifies geometric modeling in AR and supports a flexible methodology for designing in any place, smoothly transitioning from digital designs to physical outputs. By enhancing 3D creativity and accessibility through a context-aware immersive platform, we hope to advance the field of design and fabrication through AR.
2. Related Work
This section reviews the literature in three main areas: Extended reality (XR) in digital design and fabrication, in-situ tools and interfaces, and systems tailored for novice users. We provide a comparative overview of existing AR-based CAD systems, highlighting our system’s unique capabilities as summarized in Table 1.
Situated Modeling (Lau et al., 2012) MixFab (Weichel et al., 2014) RoMA (Peng et al., 2018) DesignAR (Reipschläger and Dachselt, 2019) pARam (Stemasov et al., 2024) WindowShaping (Huo et al., 2017) TinkerXR
end-to-end
in-situ CAD
built-in 3D printing
or fabrication
workflow
CAD only
CAD only
wireframe printing
CAD only
CAD only
CAD only
CSG
modeling
custom
stamping scheme
sketching
+ extrusion
sketching
+ extrusion
parametric
design
sketch
+ inflation
design
anywhere
fixed
tabletop
robotic
setup
fixed
tabletop
reference
physical objects
not
applicable
not
applicable
depth
occlusion
not
applicable
not
applicable
not
applicable
not
applicable
not
applicable
is AR headset
sufficient?
+ markers,
foot pedals, mouse
Kinect
camera setup
+ robotic
arm
+ pen,
tablet
tablet
open-source
codebase
not provided not provided not provided not provided not provided not provided
Table 1. A comparison of features of AR-based CAD interfaces in the literature.
2.1. XR in Digital Design and Fabrication
XR technologies, which include virtual reality (VR), augmented reality (AR), and mixed reality (MR), have increasingly been applied to enhance various aspects of digital design and fabrication (Dogan et al., 2022a).
Virtual Reality:
VR tools have been mostly developed for digital design compared to fabrication. Tools like Gravity Sketch allow users to create 3D models in an immersive environment, providing a spatial understanding that traditional screen-based interfaces cannot offer. Systems such as VRSketchIn (Drey et al., 2020), AdaptiBrush (Rosales et al., 2021), and HandPainter (Jiang et al., 2021) serve as interim sketching tools that can facilitate the design workflow. Researchers also studied the effectiveness of CAD tools in VR (Bourdot et al., 2010; Feeman et al., 2018; Cordeiro et al., 2019). While these tools provide immersive spatial understanding for 3D modeling, they are typically limited to virtual-only environments, separating the design process from the physical context of deployment.
Augmented Reality:
In contrast to VR, AR operates directly in physical environment, allowing users to see their designs in their intended real-world context (Iyer et al., 2025; Dogangun et al., 2024; Dogan et al., 2023c). While head-mounted displays (HMDs) can be used with hand gestures (as seen in TinkerXR) or controllers in AR (Guo and Prabhakaran, 2022; Stemasov et al., 2024; Peng et al., 2018; Jasche and Ludwig, 2020), most of the previous AR systems rely on tablets with pen or finger interactions. For instance, massless [^5] https://massless.dev/, Accessed: 19.01.2025, focuses on reducing material waste during prototyping by overlaying design information directly onto the artifact’s deployment area, enhancing precision and minimizing errors due to physical constraints. Sketched Reality (Kaimoto et al., 2022) introduces an AR sketching tool that interacts with physically actuated robots, exploring the physics of designs and mechanisms. RealitySketch (Suzuki et al., 2020) introduces a dynamic and responsive interface for sketching interactive graphics and visualizations.
TinkerXR contributes to the vision of works such as (Wang et al., 2023; Alghofaili et al., 2023; Leiva et al., 2021, 2020), which streamline digital design and fabrication workflows. However, unlike these systems, TinkerXR uniquely integrates a 3D printing workflow directly within its AR interface, providing an all-in-one solution that eliminates the need for external slicing tools or connector software while supporting precise and fully-functional designs.
Mixed Reality:
James and Eckert (James and Eckert, 2023) developed MR systems that allow operators to visualize potential outcomes projected onto the material via real-time physics simulation, enabling adjustments before a metal-cutting process begins. Other works explored conveying information spatially to users (Chen et al., 2023; Cheng et al., 2023; Li et al., 2024; Han et al., 2023). These can be integrated with digital design and fabrication systems to provide instructions or context-aware interactable menus (Dogan et al., 2024; Cheng et al., 2021; Lindlbauer et al., 2019) Window-Shaping (Huo et al., 2017) enables rapid 3D model creation on and around physical objects using a sketch-and-inflate scheme using a tablet.
Building on this foundation, TinkerXR uniquely and efficiently integrates in-situ constructive solid geometry (CSG) modeling in a seamless 3D printing workflow within an immersive AR interface. Unlike existing systems, TinkerXR allows direct 3D design and printing in any workspace without requiring additional tools or post-processing software.
2.2. In-Situ Tools and Interfaces
In-situ tools refer to systems used directly within the environment where a task takes place, rather than in separate or abstract settings. Recent advancements in AR have enabled designers to visualize changes within the actual environment, bridging the gap between artifact creation and its intended deployment (Ashbrook et al., 2016; Mahapatra et al., 2019). For example, DesignAR (Reipschläger and Dachselt, 2019) combines 2D tablet screens with stereoscopic AR and a pen for precise object tracing. RoMA synchronizes design and fabrication using a 3D printing robotic arm that incorporates existing objects into the workspace. HoloDesk (Hilliges et al., 2012) and MixFab (Weichel et al., 2014) introduce immersive tools with gestural controls tracked by Kinect, enabling context-aware designs. Unlike systems requiring fixed setups or specialized hardware, TinkerXR allows design in and for any environment using only an off-the-shelf HMD. This portability makes TinkerXR a versatile tool for diverse settings, whether at home or outside.
The concept of in-situ is not only about overlaying information but also about interacting with the environment in a meaningful way. Teachable Reality (Monteiro et al., 2023) lets users define their own in-situ tangible and gestural real-time interactions, allowing them to prototype functional AR applications without programming. SnapToReality (Nuernberger et al., 2016) enables real-time alignment of virtual objects to real-world constraints, such as edges and planar surfaces. FusePrint (Zhu et al., 2016) explored design for fitting mounts, while Robiot (Li et al., 2019) introduces designing of mechanisms that actuate other objects, and Protopiper (Agrawal et al., 2015) discusses room-scale design with low fidelity. Mitterberger et al. (Mitterberger et al., 2020; Mitterberger, 2022; Mitterberger et al., 2023) focus on larger-scale architectural designs.
Some in-situ modeling systems follow a design in and for anywhere ideology. Lau et al. introduced Modeling in Context (Lau et al., 2010), which uses everyday objects as a base point for designing around them. pARam (Stemasov et al., 2024) provides in-situ parametric design, while Situated Modeling (Lau et al., 2012) utilizes fiducial markers for robust object tracking. While these markers can impact scene aesthetics and unobtrusiveness, researchers showed it is also possible to hide them using specialized imaging techniques (Dogan et al., 2022b, 2023b). TinkerXR builds on this design in and for anywhere ideology but takes a novel direction by CSG modeling. Unlike sketching or parametric approaches, CSG modeling requires the integration of precise features such as snapping to a grid, object alignment, and robust control logic systems to enable functional and flexible design workflows. The adaptation of these traditionally desktop-bound features into an in-situ AR context is what allows TinkerXR to advance this ideology.
Tools incorporating tangible artifacts with in-situ design are valuable for context awareness and experimenting with physics, reducing the need for simulation. BrickStARt used interlocking building blocks for this purpose (Stemasov et al., 2023a). Tangible Version Control (Letter and Wolf, 2022) explores versioning of physical artifacts with the help of an HMD. Printy3D (Yung et al., 2018) teaches 3D modeling with tangible artifacts placed in a virtual container. Other tools include linkage (Jeong et al., 2018) or carving (Hattab and Taubin, 2019a, b), which make the design process interactive by touch.
Our system integrates tangibility into the design experience in several ways. It incorporates reference object types, i.e., representations of object that serve as design references, equipped with functionalities to design around them. It also features depth occlusion for physical objects, providing a realistic interaction between the real and digital worlds.
2.3. Novice Support Tools and Systems
Designing systems for novices requires consideration of simplicity, intuitiveness, and ease of learning. In the context of digital design and fabrication, this often translates into reducing the complexity of traditional CAD tools and making them more accessible. Tinkercad, for instance, has been successful in this regard by providing an entry-level CAD tool that embraces CSG modeling with simple geometric shapes and a drag-and-drop interface that is easy for beginners to grasp  (Stemasov et al., 2021). Other systems for design and fabrication using simplified core building blocks include FlatFitFab (McCrae et al., 2014), which connects planar parts for a 3D model, and similarly kyub (Baudisch et al., 2019), which uses simpler block-based design. PARTs (Hofmann et al., 2018) addresses the challenges faced by non-experts in 3D modeling by introducing a framework for expressing and reusing design intent. CraftML (Yeh and Kim, 2018) uses declarative syntax where the code dictates the output structure. Lee et al. (Lee et al., 2018, 2016) use pose and movement detection for personalized furniture design, and Tian et al. simplify analog fabrication with woodcutter  (Tian et al., 2018) and lathe (Tian et al., 2019) tools. Suzuki et al. (Suzuki et al., 2022) use virtual twins to visualize simulated behaviors when integrated with spatial references. Pfeuffer et al. (Pfeuffer et al., 2017) shows the use of virtual buttons on the wrists to access interactable menus.
Generative design has also been instrumental in achieving desired 3D outcomes with minimal effort. Some tools allow users to design and fabricate 3D models based on 2D sketches (Saul et al., 2010; Li et al., 2020; Kazi et al., 2017; Johnson et al., 2012; Igarashi et al., 1999), thereby reducing the complexity associated with directly creating models in 3D space. Other ways of generative design utilize speech inputs (Ballagas et al., 2019) and constrained parametric design (Shugrina et al., 2015). Furthermore, retrieving existing designs (Stemasov et al., 2020; Liang et al., 2022) and remixing (Follmer and Ishii, 2012; Follmer et al., 2010; Stemasov et al., 2023b; Roumen et al., 2018) enhance accessibility while balancing expressivity (Stemasov et al., 2021). ShapeFindAR (Stemasov et al., 2022) enables searching for future physical artifacts through spatial queries and text.
Among novice-oriented tools, TinkerXR takes inspiration from Tinkercad’s simplicity and accessibility, especially its use of CSG modeling. By adopting an approachable modeling paradigm like CSG, combining it with spatial interaction features, and integrating it into an end-to-end design and fabrication workflow, TinkerXR aims to lower the entry barrier for beginners and introduce new opportunities for creating designs directly in their context.
3. DESIGN RATIONALE
When developing TinkerXR, several key design considerations were prioritized to ensure the system was both effective and intuitive for novice users. The main considerations included are user-friendly design and interaction, workspace awareness, and seamless integration:
3.1. User-Friendly Design and Interaction
To enhance the user experience in TinkerXR, we prioritize both simplicity in design and intuitive interaction. Our approach combines straightforward design logic with immersive user interactions to create an accessible and engaging environment for all users. The system’s design and interaction features are crafted to ensure that even novice users can seamlessly create and manipulate models, while advanced users benefit from an efficient design process.
Simplified Design for Novices:
TinkerXR is designed with a focus on simplicity to ensure a novice-friendly user experience. By utilizing CSG operations, users can easily create complex models by combining basic shapes (e.g., cuboids and spheres) through simple boolean operations such as union and subtraction. This approach reduces complexity, allowing users to design without managing technical parameters. The system categorizes shapes into solid and hole types, enabling the addition or removal of material to create more advanced designs effortlessly.
Intuitive Interaction:
Interactivity in TinkerXR is driven by hand gestures and movement, recognized by Quest’s tracking system. These gestures allow users to manipulate objects directly, making the design process intuitive and engaging. Building upon the methods outlined in prior works (Kim et al., 2005; Weichel et al., 2014; Stemasov et al., 2024), TinkerXR lets users interact with the elements directly with their hands and without the need for external tools other than the headset. For instance, a pinching motion allows users to grab and move objects, while wrist rotation toggles features such as snap-to-grid, multi-selection, and uniform scaling. Moving the weak hand along a single global axis locks object manipulation to that axis, ensuring precision and control. Wrist-mounted menus provide an intuitive interface with buttons for quick navigation, including shortcuts for pinch gestures using fingers other than the index (e.g., select/deselect all). Visual and auditory feedback confirms actions, further enhancing the system’s usability and creating an immersive experience.
3.2. Workspace Awareness
TinkerXR’s real-world integration is achieved through sophisticated scene recognition and depth occlusion technologies. These allow the system to not only understand the environment but also intelligently interact with it.
Figure 2. Real-life workspace awareness. (a) Three cases for selecting the workspace in a scanned room. (b) Grid appearing on the selected workspace plane for the three corresponding cases. (c) Selecting the workspace based on a reference of a physical object. (d) Physical and digital objects rendered together with depth occlusion off (upper) versus depth occlusion on (lower).
[…]
Four photos labeled (a), (b), and two in (c). The first two photos show the first-person view of the user looking at a room with a table inside. The other two photos show three small decorative cacti on a table with a digital pyramid model behind them. In the image (a), three yellow rays come out from the user’s right hand, first hitting the floor where a text appearing above says ”select: floor”, second hitting the top of the table where a text appearing above says ”select: table”, and the third hitting the wall where a text appearing above says ”select: wall”. There is a white outline around the table. In image (b) the room is seen from the same angle as an image (a), but this time there are grids on top of the floor, table, and wall. In the first image of (c), the digital pyramid model partially blocks the view of the cacti even though it is placed behind them while in the second image of (c) the cacti are fully visible.
Scene Recognition:
It is essential for the system to recognize and adapt to its physical surroundings by using real-time environment mapping to adjust designs according to available space and nearby objects. The system can identify furniture and other objects, categorizing them as potential design spaces or obstacles, such as a desk, wall, or floor. This recognition ensures accurate placement and scaling of models within the environment. For instance, the system’s awareness of the selected workspace allows it to overlay a grid layout, enabling users to utilize snap-to-grid features and visualize object projections on the workspace as a reference during the design process. Additionally, the system supports the creation of reference objects, which are simple wireframe geometries representing physical items (e.g., a plant pot when designing a plant pot holder). These reference objects enable users to place the workspace grid plane on them, facilitating designs that seamlessly integrate with real-world items.
Depth Occlusion:
It is also essential for users to understand their surroundings for context awareness while designing. While AR functionalities significantly aid in this area, models rendered in front of every world object can detract from the realistic nature we aim to achieve in the design task. By understanding the depth of objects in the environment, TinkerXR can occlude parts of a model that are behind physical objects, enhancing the realism and immersion of the design experience while elevating the user’s perception of the design as part of their real environment.
3.3. Seamless Integration
The transition between designing in AR and preparing a model for fabrication (in our case, 3D printing) should be fluid and straightforward, encouraging users to iterate on designs quickly and efficiently.
Integrated Design and Fabrication:
In TinkerXR, we have consolidated design and fabrication functionalities into a single system, streamlining the entire workflow. Users can design their models and select printing options (e.g., layer height, infill pattern, support structure) directly within the AR environment, eliminating the need to transfer designs between separate tools. This integration simplifies the process by allowing users to handle all aspects of design and preparation within a unified interface, thereby reducing the complexity traditionally associated with using multiple software packages.
Direct Communication:
Our system further enhances the user experience by enabling direct communication with WiFi-enabled 3D printers from within the system. This capability eliminates the need for a separate computing device and removes intermediate steps, such as exporting to slicing software, by sending designs directly to the printer.
4. Implementation
Figure 3. User workflow. (a) Users begin by selecting a workspace from one of the flat surfaces in their room. (b, c) On the selected surface, users can create and manipulate primitive geometries using their hands while utilizing the tools provided by TinkerXR. In this example, the user designs a finger skateboard obstacle. (d) The final model is then placed in a virtual 3D printer twin where print slicing options can be adjusted before (e) fabricating on the remote physical printer. (f) Examples from the user study (Section 6) that demonstrate the design anywhere functionality, i.e., the flexibility to adapt to different environmental contexts.
[…]
TinkerXR was implemented for Meta Quest 3. The system code is written in C# with Unity Version 2022.3.22f1 LTS [^6] https://unity.com/releases/2022-lts, Accessed: 19.07.2024. The system utilizes open-source resources provided by Meta to support various features. Specifically, we used Meta XR All-in-One SDK [^7] https://developer.oculus.com/downloads/package/meta-xr-sdk-all-in-one-upm/, Accessed: 19.07.2024, which includes libraries for hand tracking, room and furniture detection (Scene API [^8] https://developer.oculus.com/documentation/native/android/mobile-scene-api-ref/, Accessed: 2.08.2024), passthrough features, and other essential XR development tools for Unity. Additionally, we incorporated Meta’s Depth API [^9] https://github.com/oculus-samples/Unity-DepthAPI, Accessed: 19.07.2024 library for the depth occlusion feature and Unity’s built-in audio engine for the spatial audio utilities. The subsequent sections will outline the envisioned procedural steps of using TinkerXR with its features, from (1) selecting the workspace to (2) designing artifacts, and finally to (3) fabrication of those artifacts.
4.1. Workspace Selection and Modification
TinkerXR system enables users to select their desired workspace, which can vary from the top of a desk to the floor, walls, door frames, and other flat surfaces.
4.1.1. Scene Setup
First, users scan their rooms using the headset’s built-in space setup feature. This feature identifies the room’s boundaries and furniture while providing users the option to manually configure their furniture (e.g., for selecting which part of the desk they would like to work on). The scanned furniture is saved as rectangular prism models, representing all selectable workspaces as simple, flat surfaces (Figure 2a).
4.1.2. Workspace Selection
After scanning, users can interact in TinkerXR and view the highlighted selectable workspaces. This is possible with the headset’s built-in room recognition feature and MRUK framework [^10] https://developer.oculus.com/documentation/unity/unity-mr-utility-kit-overview/, Accessed: 20.07.2024. Then, users use their dominant hand to direct a ray toward the surfaces of furniture or room boundaries, which displays the name of the planned workspace it intersects (Figure 2a). An index finger pinch gesture is then employed to confirm the selection.
4.1.3. Workspace Modification
After the workspace is selected, a grid is overlaid on the chosen surface, covering the entire area of the furniture’s selected side or the boundary (Figure 2b). This grid facilitates snap-to-grid operations, enhancing precision in the design process. While this grid is visually planar, its snap points extend alongside the grid plane’s normal direction, spreading snap points into the scene space in all dimensions. The density of the lines on the grid depends on the grid spacing and by default, the grid spacing is set to 2 cm, with the option to adjust this value via the hand menus. Additional grid options include selecting its color, which enhances visibility (e.g., a white grid on a white surface may be less visible), determining whether physical objects should occlude the grid, and adjusting its distance to the surface alongside its normal, which can be useful against z-fighting issues when the grid is in occlusion mode. Additionally, depth occlusion for digital objects can be toggled on and off (Figure 2c), aiding in designing with physical world constraints and designing artifacts that interact with real-world objects.
4.2. Design Features
Once users select their workspace, they can begin designing models directly with their hands, leveraging an array of tools typically found in traditional CAD systems.
4.2.1. Object Creation
Users can choose from seven types of primitive 3D geometries (cube, sphere, cylinder, capsule, triangular prism, pyramid, and cone) using buttons on their weak hand wrist (Figure 4a). The color of the button turns blue when the finger hovers close to the button. This gives the user feedback on which button is about to be pressed. Upon pressing a button, the selected object appears on the user’s weak hand (Figure 4b and 4c).
Figure 4. Object creation and selection. (a, b, c) Poking gesture on buttons for creating primitive objects. (d) Highlighting by pointing hand towards an object and selecting it by pinch move in single selection mode. (e) Selecting objects in multiple selection mode.
[…]
Five photos labeled (a), (b), (c), (d), (e) show the first-person view of a user in front of a wall interacting with their hand and the digital objects. In images (a), (b), and (c) user have extended their weak hand to their eye level and looking at the inside of their wrist. There are 12 white square buttons on the wrist in a three-by-four grid. Seven of the buttons show primitive geometric objects, namely cube, cylinder, sphere, capsule, triangular prism, pyramid, and cone. In the image (a), the user is hovering his dominant hand index finger on their wrist. In the image (b), the user pokes the cube button with their dominant hand index finger and the button turns green and a cube appears above the weak hand. In the image (c), the user does the same with a sphere next to the cube. In image (d), user does a pinch motion while their dominant hand hand is directed toward the created cube and their weak hand is oriented palm-down. The cube’s color is changed from blue to cyan and a white grid is appearing on the cube. Similarly, in the image (e), the user does the same pinch motion for the sphere, this time with the weak hand oriented palm-up. Both of the objects turn to cyan, with white grid appearing on the sphere instead of the cube.
4.2.2. Object Highlighting and Selection
During interactions, users may encounter multiple instances of objects, necessitating methods for selecting and manipulating specific items.
Highlighting
To select an object, users point their hand toward it, with an invisible ray cast from the tracked hand model in the direction of the pointing finger. When this ray intersects with an object, visual feedback is provided (Figure 4d). Highlighted objects are rendered with a white grid overlay featuring transparent lines, ensuring the object’s original color remains visible.
Selection
A pinch gesture, i.e., thumb and index finger touching, selects a highlighted object (Figure 4d), turning it bright cyan and emitting a directional sound. A surrounding Object Manipulation Box appears, enabling movement, resizing, and rotation.
4.2.3. Single and Multiple Selection
Selection modes toggle based on weak-hand orientation: palm down for single selection (default, Figure 4d) and palm up for multiple selection (Figure 4e).
Single Selection
In the single selection mode, only one object is active; selecting another deselects the current object. As this is considered the default mode, the palm-down orientation is chosen as it is the resting position of the hand.
Multiple Selection
This mode enables group manipulation, i.e., the equivalent of pressing the control key on a keyboard in desktop interfaces. Selecting an object that is not already selected adds it to the list of selected objects. Conversely, selecting an already selected object removes it from the list. Multiple selected objects can be manipulated together at once.
4.2.4. Object Manipulation
The Object Manipulation Box facilitates object movement, resizing, and rotation.
Figure 5. Moving, rotating, resizing objects. (a) Moving objects around the world space with the movement handle. (b) Rotating objects by twisting the rotation handle around its corresponding axis. (c) Resizing objects with the corner handles.
[…]
Six photos labeled (a), (b), (c) each label corresponding to two photos show a cube being manipulated with their dominant hand from the first-person view of the user in front of a wall. The first pair of pictures labeled (a) shows the user pinching the movement handle on top of the cube in the first picture and dragging their dominant hand and the object to the right in the second picture. The second pair of pictures labeled (b) shows the user pinching the rotation handle next to the cube in the first picture and rotating their dominant hand wrist and the cube in the second picture. The third pair of pictures labeled (c) shows the user pinching the corner handle at the top right corner of the cube in the first picture and moving their hand to the right and the cube stretching to the right in the second picture.
Moving Objects
A movement handle, i.e., gizmo, above the objects follows the user’s pinch gesture, which allows for freehand movement (Figure 5a).
Rotating Objects
Three axis-aligned rotation handles allow rotation around the x, y, and z axes. Grasping a handle with a pinch gesture rotates objects like a knob (Figure 5b).
Resizing Objects
Corner handles at each vertex of the Object Manipulation Box enable resizing. Adjusting a handle moves it freely while the opposite handle remains fixed, which ensures precise control (Figure 5c).
Uniform scaling activates when the weak hand is sideways (Figure 6a, 6b). A guiding line extends from the box center through the selected corner, ensuring accurate scaling along this line. Multiple selections allow synchronized scaling, while free scaling does not preserve aspect ratios (Figure 6c).
Figure 6. Uniform vs. non-uniform scaling. (a, b) Resizing objects with uniform scaling mode. (c) In free scaling mode.
[…]
Three photos labeled (a), (b), and (c) show a first-person view of a user looking at a wall and interacting with a complex 3D abstract design. In images (a) and (b), the user has their weak hand’s palm looking right and they apply a pinching motion to the corner handle on the upper right side of the manipulation box. A pink ray is seen coming from the middle and going through the selected corner handle. The user brings their right hand in the second picture closer to the center of the manipulation box, resulting in a smaller object with the same aspect ratio. Image (c) shows the user dragging the same corner handle with their weak hand palm looking down. The resulting object is compressed vertically.
Snap-to-Grid
When any of the elements for manipulating objects are held with the dominant hand with a pinch gesture, the orientation of the weak hand dictates the toggle between free movement and snap-to-grid modes instead of single and multiple selection modes. Free movement is enabled when the weak hand’s palm faces upward, while snap-to-grid is activated when the palm faces downward. The functionality adapts to movement, rotation, and scaling operations.
Figure 7. Manipulating objects with snap-to-grid mode for (a) movement, (b) rotation, and (c) resizing.
[…]
Three photos labeled (a), (b), and (c) show the first-person view of the user looking at a wall with a vertical grid on the wall, holding their weak hand palm looking up, and manipulating a cube object. In image (a) user does a pinch motion to the movement handle on top of the cube. There are arrows appearing perpendicular to the sides of the cube with uniform steps. Similarly, in image (c), the same arrows are seen but for the scaling handle on the corner of the cube. In image (b), the user applies a pinch motion to the rotation handle on top of the cube and a circle with 15 degrees of steps parallel to the top side of the cube is seen.
For movement, when the snap-to-grid feature is activated, the Object Manipulation Box will align with the grid lines on the selected workspace (Figure 7a). In some cases, both parallel sides of an object may not be able to snap to the grid lines due to the distance between them. In such situations, the system determines which side should snap to a grid line based on the following logic:
  1. (1)
    If a side is parallel to the grid plane, the system will always favor the side closer to the grid. (e.g., placing the bottom of the object on the table grid)
  2. (2)
    If a side is not parallel to the grid plane, the system will favor the side that aligns with the direction of movement calculated from the start location of the grabbing action.
For rotation, when the snapping option is activated, the rotation handle adheres to a 15-degree increment from its initial position (Figure 7b). This 15-degree step is chosen to provide a balance between precision and ease of use, ensuring that users can achieve accurate rotations without making the process cumbersome.
For scaling, when the snap-to-grid feature is activated, the corner handle aligns with the nearest grid point within the scene space to the pinch location (Figure 7c). The objects are scaled in the same logic as the free movement mode, where the opposite corner acts as an anchor.
4.2.5. Axis Locking
The axis lock tool in TinkerXR allows users to easily constrain object manipulation to specific axes or planes, enhancing precision during design. Users can activate the axis lock tool by performing an index finger pinch gesture with their weak hand. Moving their hand in the desired global direction triggers a visual cue in the form of a colored edge (Figure 8a) or a side (if moved diagonally) of the axis lock tool, representing the active axis or plane of manipulation. This interaction ensures that users can limit movement or resizing action along a single axis or within a plane (Figure 8b and 8c) without needing to interact with additional buttons or menus. A single tap using the pinch gesture with their weak hand resets the axis lock, allowing users to resume unconstrained manipulation.
Figure 8. Axis locking. Users can (a) initiate the axis lock by pinching with their weak hand and moving it in the desired direction. Useful for limiting the (b) movement and (c) resizing of the objects.
[…]
4.2.6. Solid and Hole Objects
Users can toggle whether an object is solid, adding material, or hole, removing material, by pressing a button on their weak hand wrist. The hole objects appear with a transparent mask, making them identifiable when compared to the solid objects (Figure 9a). The hole objects can be manipulated the same way as the solid objects.
4.2.7. Combining Objects
Users can combine and turn selected objects into one uniform object with a button on the weak hand wrist. The combined object will turn into the color of the last solid object chosen. The combining operation works with solid and hole objects, removing the hole objects after performing a cut on the solid objects through the intersection between the solid and hole objects (Figure 9b and 9c).
Figure 9. Multi-object operations. Users can create (a) solid and hole objects. (b, c) The combining operation creates a cut in the intersection between the solid and hole object.
[…]
Three photos labeled (a), (b), and (c) show a first-person view of a user looking at a wall. In the image (a) a solid cube and a hole sphere intersecting with the top corner of the cube is seen. The hole (or negative) sphere is seen with red color and with dotted transparent skin. In image (b) the weak hand wrist is visible while the cube and the sphere are selected and colored cyan. The user is about to poke a button named ”combine” with their dominant-hand index finger. In the image (c), the combination of the objects is visible where there is a round cut on the upper corner of the cube where the sphere used to intersect the cube.
4.2.8. Reference Objects
TinkerXR supports creating reference objects, offering users a powerful tool for designing around physical items. Users can activate the reference mode via the hand menu on their dominant hand wrist (Figure 10a), which converts default block objects into reference objects. These reference objects are transparent and feature highlighted wireframe edges, distinguishing them from standard objects.
This wireframe structure allows users to precisely position grids on reference objects, aligning them with vertices and plane normals for accurate design. This translates to the grid not only functioning horizontally or vertically (as seen on the common workspaces) but also diagonally in any direction, fitting to the selected surface of the reference object. The transparency aids in manipulating these objects to replicate physical counterparts (Figure 10b,c), enabling users to build around them easily. Additionally, reference objects can be set as hole types, letting users capture intricate details of physical objects for better integration.
Figure 10. Reference objects. Users can (a) create reference objects via the dominant-hand menu and (b) use them to represent physical items. (c) On-object grids are visualized for precise design. (d) The example shows a plant pot hanger designed around a reference pot.
[…]
4.2.9. Ruler Tool
Users can create a virtual ruler by pressing a button on their weak hand wrist. This simple ruler, equipped with two interactable edge handles (similar to scaling handles), provides real-world scale measurements. This ruler can be used for precise alignment and measuring distances between objects within the real and virtual space (Figure 11a).
Figure 11. In-context measurement. Users can (a) measure the length of the objects around them with the ruler tool. For parametric resizing, users (b) select the handles on the edges of the object manipulation box, which (c) activates a numerical keyboard on the weak hand for (d) inputting the desired length alongside the edges.
[…]
Four images labeled (a), (b), (c), and (d) show a user’s first-person view using the system. In picture (a), user is using their weak and dominant hand to pinch both ends of a digital line where in the middle of the line it says ”20.00 cm”. The user is measuring out the length of a game console (Nintendo Switch). In pictures (b), (c), and (d) user is manipulating a digital cube model. In the picture (b) user is applying a pinch motion to a small black handle on the middle of the edge of the cube. In picture (c), a numerical keyboard interface appears on the weak hand in direction of the palm and it says ”10 cm” in the input field of the keyboard interface. In picture (d), the user has inputted ”20 cm” as the input and clicks the confirmation button. The cube’s width is doubled.
4.2.10. Parametric Resizing
In addition to free, snap-to-grid, and uniform scaling operations, users can also adjust object sizes through precise numerical length inputs. To use this feature, users first select the edge of the Object Manipulation Box they wish to resize (Figure 11b) using a pinch gesture. Then, a numerical keyboard appears on the user’s weak hand (Figure 11c), allowing them to enter the desired real-world length for the selected edge. The object scales proportionally along the direction of the chosen edge based on this input (Figure 11d). This method can be effectively combined with the ruler tool to ensure accurate design according to specific measurements.
4.2.11. Other Functions
Other buttons placed on the dominant hand wrist enhance the efficiency of the design process. These features include selecting the color and material of the objects, duplicating selected objects, selecting or deselecting all objects, deleting selected objects, deleting all objects, undo, redo, saving, and loading operations.
To streamline the selection process, pinch gesture shortcuts were added. A pinch between the thumb and the middle finger triggers the function to deselect all objects, while a pinch between the thumb and the ring finger triggers the function to select all objects.
4.3. Fabrication Features
After the target model has been created, the user can export it to an STL file. This will allow users to access it from other CAD tools and enable them to print their designs. This feature not only accelerates and simplifies the fabrication process but also enables users to quickly fabricate a draft of their design to assess whether it will be printed with the fidelity they anticipate.
The implementation of the server system was possible with the usage of Docker containers to package the working environment for different computers. The server itself was written with JavaScript in Node.js, and the slicer used for the server was Ultimaker CuraEngine [^11] https://ultimaker.com/software/ultimaker-cura/.
Figure 12. Transition to fabrication. Virtual printer twin is used as a placeholder for exporting the model for a target printer or conducting a quick size check. Users can (a, b) drag and drop the models they create in the printer, (c) modify the print orientation, and (d) instantiate physical printing.
[…]
Four images labeled (a), (b), (c), and (d) show a user’s first-person view looking on top of a table where there is a horizontal grid appearing on the table and a virtual 3D printer twin and a model of a wall hook is visible. The 3D printer twin is seen as a transparent rectangular prism with white edges and a grid on the bottom plate. The name of the printer (Creality Ender-3 V3 KE) is written on the bottom of the printer. In the picture (a), the user is moving the model with the movement handle grabbed with a pinch gesture with the dominant hand, dragging it to the printer twin. A white line grid appears on the printer twin. In picture (b), the model is placed in the middle of the printer twin and it says ”model printed successfully” in green color on top of the printer twin. On the right side of the printer twin, two input fields and two buttons are visible. The first input field is labeled ”enter sever IP…” and the second input field is labeled ”enter printer IP…”. The first button is labeled ”Start print” and the second button is labeled ”Slicer settings”. In the picture (c), user is rotating the model with a pinch gesture applied to the rotation handle on top of the model. In picture (d), a progress bar is seen on top of the printer twin and a text is written as ”0%” in the middle. The input fields contain the partially blurred IP addresses of the server and the printer and the buttons are labeled ”pause print” and ”abort print”.
Figure 13. Preparing for printing and slicing. (a) Users can select the slicing options for their print. They initiate slicing (b) and (c) view the sliced preview of the model, with outer walls in gray and support structures in cyan. (d) If a model exceeds the printer’s build area, a ”too large for printer” indicator appears above the printer twin.
[…]
The low-poly Stanford bunny is placed inside the virtual printer twin with slicing options displayed next to it.
4.3.1. Spatial Virtual Twin of Printer
A printer twin is a unique object that cannot be selected in the multiple selection mode or cannot be merged with other objects. This virtual printer object helps the user estimate how the object will be printed inside the real-world 3D printer. The size of the virtual printer twin corresponds to the real-world printer’s build area. This virtual printer’s size can be selected from the predefined printers or set manually. To start the printing process, the user has to drag and drop the current model into the printer twin. If neither the server IP nor printer IP is given on the printer twin, it will save the current model locally into an STL file. If the slicer server is connected to the headset, the user can send their STL files to the server to create a G-code file for 3D printers. If the printer is also connected to the server, then the printing process will begin automatically.
4.3.2. Drag and Drop and Adjustments on Build Plane
Users can pick up the models they created, and drag them to the virtual printer twin with the movement handle, where the printer will be highlighted (Figure 12a). After a model is dropped over the printer twin, the duplicate of the objects comprised by the model will be placed inside the printer (Figure 12b). The original objects will be moved to their initial position. After the object is placed inside the printer, users can adjust the size, orientation, and position of the object (Figure 12c). Note that a print can not be initiated when the objects go outside the boundaries of the build area.
4.3.3. Preparing for Printing and Slicing
The sliced model is visualized in the virtual printer (Figure 13b and 13c), where the main walls are displayed in gray and support structures in cyan (tree support selected). The visualization allows users to review the slicing results and identify areas requiring adjustment before starting the print. If the selected model exceeds the printer’s build area, the system notifies users by displaying a ”too large for printer” indicator above the virtual printer twin (Figure 13d). This ensures that users can modify their design or scale it appropriately before initiating the print.
After the printing request is sent to the virtual printer twin object, if no server IP is given, the STL model of the object in the virtual printer will be saved locally. If the server IP is given while the printer IP is not, this STL model and sliced G-code file will be saved on the remote server. However, if both IPs are given, the printing process will start after slicing. The headset monitors those steps. The server IP is the same as the computer the server program is running on, and the printer IP can be obtained from the printer after it has been started and connected to the internet. Note that the printer should be accessible from the server computer if they do not reside on the same subnet.
The server manages multiple slicers to serve the XR headset clients. An example slicer configuration for testing has been created for Ultimaker S3121212https://ultimaker.com/3d-printers/s-series/ultimaker-s3/ printer with options: Enable Support, Support Structure, Layer Height, Infill Density, and Infill Pattern (Figure 13a). More slicers can be implemented by adding them to the slicer_options.json file. The slicer options are completely modifiable and could be increased with ease as the run command is also read from the JSON. How the slicer should work is completely in the hands of the programmer as the command could be easily manipulated by the flexible replacement tags that have been noted as options. The only required arguments that a slicer should take are output and input file arguments, which use gcode_path and stl_path, respectively.
A client can make multiple requests on the server. When a /slice GET request is sent to the server, it returns the available slicers and their options. On the other hand, when a /slice POST request is sent to the server, the STL file is loaded on the appropriate slicer, and the G-code file of the target model is obtained. The server parses the options selected by the user and passes them to the slicer. After the slicing operation has been completed, the resulting G-code file is visualized in the virtual printer. This draft object is shown in different colors for the main and the support parts. After that, the printing process can start. Note that one server can handle multiple printers, and one headset can connect to various servers. The only necessary action is to enter IP addresses.
4.3.4. Printing and Monitoring
After the /print POST request is sent to the virtual printer twin object, the server will connect and get access permission to the 3D printer. This permission is only necessary the first time as the server will save the token for accessing the printer, and access permission will not be necessary until the token has expired. If access to the 3D printer is granted, the server will send a printing request with the G-code file to the printer. The headset can obtain the print process information by sending the /print GET request. The user can pause or abort the printing process by sending a PUT request to the /print URL if necessary. This request supports three different commands: continue, pause, and stop, which can be sent via the buttons next to the printer twin (Figure 12d).
5. Applications
In this section, we present ten distinct designs created with TinkerXR (Figure 14) to demonstrate its versatility and capabilities across various workspace environments. These examples highlight how TinkerXR enables users to easily produce both practical and creative objects through spatial interaction, tailored to their immediate surroundings.
Figure 14. Designs created via TinkerXR on the left; their fabricated results on the right, placed in their intended space. These are (a) phone holder, (b) coaster, (c) pen holder, (d) candy bowl, (e) finger skateboard obstacle, (f) figurine stand, (g) key hanger, (h) door wedge, (i) plant pot hanger, (j) lamp shade.
[…]
Ten images labeled (a), (b), (c), (d), (e), (f), (g), (h), (i), and (j) show ten different 3D printed results next to their virtual rendered versions. The image (a) has a phone holder with a smartphone on top. In the image (b), there is a drink coaster with a cup on top of it. The image (c) has a pen holder with colorful markers inside it. In the image (d), there is a candy bowl with packaged candy inside. In the image (e), there is a ramp and a grinding rail with a finger skateboard next to it. In image (f), a figurine stand is mounted to a wall with a small figurine. In the image (g), there is a wall key hanger with two hooks and a bunch of keys hanging down those hooks. In the image (h), there is a door wedge holding a door from closing. In image (i), there is a plant pot carrying a ring connected to ropes that are hanging from the ceiling. In the image (j), there is a spherical light bulb shade covering a light bulb mounted to the ceiling.
5.1. Desk Workspace
In the desk workspace, five designs were developed: a phone stand, drink coaster, pen holder, candy bowl, and fidget terrain (Figure 14a-e). These objects demonstrate the system’s ability to facilitate personalized designs suited for everyday use. The phone stand and drink coaster highlight basic cutting operations and represent simple designs in everyday objects. The pen holder showcases intricate decorative patterns, made by combining and duplicating hole objects in a chain, and the candy bowl explores experimental shapes by combining rectangular prisms into curvilinear forms. Fidget terrain stands out as a playful interactive object designed with ramps and rails, which illustrates the potential for creating designs that interact with moving physical objects. Each design benefits from real-time in-situ customization, ensuring the objects fit and function precisely within the user’s workspace. Here, working on smaller scales with hand interactions can be challenging; thus, a useful approach is to create the design on a larger scale and then scale it down to the intended size. This helps manage precision and detail when working with smaller components.
5.2. Wall Workspace
For the wall workspace, two objects were designed: a figurine stand and a key hanger (Figure 14f and 14g). The figurine stand allows users to display small decorative items without adding clutter, created using the ruler tool for a fitting hole. The key hanger offers practical functionality with aligned hooks using snap-to-grid. The models are snapped to the wall while designing for a better sense of depth, helping users create visually integrated solutions.
5.3. Other Workspaces
Beyond the more common desk and wall spaces, TinkerXR can also be applied to other environments, e.g., ground, ceiling, doors, windows, screens, couches, and more. On the ground, a door wedge was designed to hold doors securely in place (Figure 14h), benefiting from real-time scaling for accurate size.
For the ceiling, we created two distinct designs: a plant pot hanger and a light shade (Figure 14i and 14j). The plant pot hanger demonstrates the system’s ability to accommodate objects that require support and suspension, such as using ”ropes” for hanging, a practical solution for overhead applications. The light shade, on the other hand, highlights the system’s capability to handle complex and abstract 3D modeling. Its artistic design illustrates the system’s flexibility in creating intricate structures on a headset without technical limitations. These examples show the adaptability to less conventional design spaces and orientations.
6. Evaluation
To evaluate TinkerXR, we conducted a user study involving three design tasks of varying contexts, comparing to Tinkercad as a baseline. We used both quantitative measures and qualitative feedback to gauge user performance.
6.1. Methods
Participants. We recruited ten participants (ages 23-29; 9 male, 1 female). Four participants had little experience with CAD software, and six had no previous CAD experience. Additionally, three participants had little experience with AR systems, and seven had no experience with AR systems. Finally, four have used 3D printers in some parts of their life, while six have never used a 3D printer.
Figure 15. Designs created by the participants in the user study. Objects are colored based on their task number (a: wall-mounted key hanger, b: cable management, c: toiletries holder) and the system used for designing the models.
[…]
Task & Procedure. For the evaluation, we conducted a within-subjects comparative study to assess TinkerXR against Tinkercad. Each participant completed modeling tasks using both tools, with the order of tool usage counterbalanced across participants to mitigate order effects (e.g., half of the participants used TinkerXR first, followed by Tinkercad, while the other half did the opposite). We showed both systems briefly, and participants completed a 3-minute warm-up before starting the tasks to familiarize themselves with the tools. Participants were assigned three modeling tasks, each tailored to a specific context and accompanied by explicit design requirements: #1 Wall-mounted key hanger: Participants created a key hanger with two hooks narrow enough to accommodate keychain rings. This task emphasized designing small, precise features and required the use of a vertical wall as the design context. #2 Cable management: Participants designed a holder to organize cables on a desk. The design had to securely fit the provided cables, ensuring they would not fall, and the organizer needed to remain stable when attached to the desk. #3 Toiletries holder: Participants designed a holder for a toothpaste tube and toothbrush, intended for use in a bathroom. They were provided with physical objects representing the toiletries and used the reference object functionality to ensure the holder dimensions aligned with the objects. For both conditions, participants were allowed to use a physical ruler to measure physical dimensions, although they were also allowed to use the digital AR ruler in the TinkerXR condition.
Participants engaged with TinkerXR using a Quest 3 headset and hand interactions. For TinkerXR, they were instructed to prepare and start their prints on a virtual twin of the 3D printer; while for the Tinkercad condition, they used Cura [^13] https://ultimaker.com/software/ultimaker-cura/’s desktop UI for 3D printing preparation. To save time, participants did not wait for the physical prints to finish before moving on to the next task. The printer slicing sequence was only carried out for the first task due to the same reason. Each modeling session was followed by a survey to gather qualitative and quantitative feedback on the tools, focusing on usability, efficiency, and overall satisfaction.
Measures. We used both quantitative and qualitative metrics to evaluate the system. We recorded the time required to complete the individual 3D design tasks as a measure of overall system performance. For Task 1, we also measured the time required to transfer the finished design into the 3D printing environment and prepare it for fabrication, i.e., slicing. We measured this duration until the participant hit the print button. After completing the tasks, participants were asked to complete standardized questionnaires, including the System Usability Scale (SUS) (Brooke, 1996), which we extended with additional questions related to the immersive experience, and the NASA Task Load Index (NASA-TLX) (Hart and Staveland, 1988). We collected qualitative feedback through open-ended questions regarding their overall experience.
6.2. Results
6.2.1. Observations and Analysis of the Fabricated Artifacts
Figure 15 shows the three types of objects designed and fabricated by the participants. Each design task had specific requirements to ensure functionality and usability, and all participants successfully met these requirements. For the key hanger, the design required at least two hooks to hold keychains. The hooks had to be narrow enough to accommodate standard keychain rings. For the cable management, participants needed to design a holder that securely fit at least two cables of provided diameters, ensuring the cables would not slip out. Additionally, the design had to accommodate the thickness of the desk to remain stable when attached. For the toiletries holder, the design required at least two compartments or holes to hold the provided toothbrush and toothpaste tube. These compartments had to match the given dimensions to ensure a stable hold. All models were validated by checking if they adhered to dimensional requirements and functionality constraints defined for each task.
A notable insight was that cube primitive as a solid and cylinder geometry as a hole were mostly used objects for both the TinkerXR and Tinkercad conditions. During TinkerXR usage, snap-to-grid was heavily utilized. All of the participants used the physical ruler for the wall-mounted key hanger and cable management tasks while designing with Tinkercad, while all of them used the reference object functionality of TinkerXR for the toiletries holder while designing with TinkerXR. Participants relied heavily on depth occlusion and the highlighted projection of the objects on the grid to understand the location and depth of the objects. Seen in Figure 15, the outcomes that have taken the most time on average of two systems were the key hanger by P5 (18m 45s), the toiletiers holder by P5 (15m 31s), and the key hanger by P9 (14m 35s).
All of the resulting objects with TinkerXR vs. Tinkercad were comparable, i.e., none of the users created designs that differed deeply depending on the tool used. This verifies that TinkerXR was able to offer similar 3D CAD capabilities despite the newly introduced immersive AR environment and not being an established commercial solution.
6.2.2. Time
Across all three design tasks, we found no statistically significant difference in task completion time using TinkerXR compared to Tinkercad: Task 1 (t = 0.272, p = 0.792), Task 2 (t = 0.272, p = 0.792), Task 3 (t = -1.769, p = 0.110). This suggests that TinkerXR provides an immersive design experience without a substantial time cost. However, future work may explore this tradeoff in larger or more varied user populations, e.g., for experts as well. Fig. 16 shows the results for all tasks. When it comes to the slicing process for Task 1, the 3D printing preparation took less time when participants utilized the TinkerXR system. A paired t-test confirmed significantly shorter completion times in the TinkerXR condition for Task 1 (t = 3.747, p = 0.004).
Figure 16. Recorded task completion times (in seconds) for participants using Tinkercad and TinkerXR across three design tasks and one slicing task. No statistically significant difference in completion times was found between the two systems for the design tasks. For the slicing task, participants completed the process significantly faster with TinkerXR compared to Tinkercad (t = 3.747, p = 0.004).
[…]
Result of the recorded times of the participants.
6.2.3. Surveys
For the standard SUS survey, we obtained a score of 90%, which shows that TinkerXR system is within the A grade (90% or above), while Tinkercad achieved a score of 85%, which places it within the B grade (from 80% to 90%)  (Sauro and Lewis, 2016). As shown in Fig. 17, the TinkerXR system has superior scores across statements such as Q3, Q9, and Q10. These findings support that TinkerXR was easier to use for users in terms of coherency. The additional questions related to immersive experience show that TinkerXR helped users to have a more natural and hands-on designing procedure.
Figure 17. Results of the extended SUS survey comparing Tinkercad and TinkerXR. User ratings are shown on a Likert scale from 1 (strongly disagree) to 5 (strongly agree). TinkerXR demonstrates generally higher ratings for immersive design features.
[…]
Results of the SUS survey.
The analysis of the NASA-TLX survey with a paired t-test reveals significant differences between Tinkercad and TinkerXR across multiple workload dimensions. Participants reported lower mental demand when using TinkerXR compared to Tinkercad (t = 3.857, p = 0.004), indicating that the immersive AR environment facilitated easier comprehension and interaction during design tasks. In contrast, TinkerXR exhibited higher physical demand (t = -5.658, p <0.001) and effort (t = -2.684, p = 0.025), likely due to the reliance on hand gestures and the physical interaction required with the AR system compared to the mouse-and-keyboard interface of Tinkercad. However, temporal demand, performance, and frustration levels were comparable between the two tools, indicating that the usability benefits of TinkerXR may mitigate the added physical strain.
Additionally, considering the current capabilities and limitations, users indicated that they would use TinkerXR for future projects, with an average rating of 4.4 on a 5-point Likert scale. They also recommended this tool to other novice users, giving it an average rating of 4.7 out of 5.
Figure 18. Results of the NASA-TLX survey comparing Tinkercad and TinkerXR across six workload dimensions. TinkerXR shows lower mental demands and higher perceived performance, while Tinkercad exhibits lower physical demand. Performance, temporal, and frustration scores are comparable between the two tools.
[…]
Result of the NASA-TLX survey.
6.3. Qualitative Feedback
Open-ended questions in our survey, such as ”What did you like most about using the tool?”, ”What challenges did you face while using the tool?”, and ”What aspects of the tool would you most like to see improved?”, and the thematic analysis of the responses allowed us to gather deeper insights into the system’s strengths and areas for improvement under five overarching themes:
Real-World Integration and Immersive Experience
Participants consistently praised the ability to design within real environments, emphasizing the value of TinkerXR’s AR capabilities in visualizing objects in situ. For instance, P1 shared, “With TinkerXR, I could see how my design would actually look in the space it was meant for.” P4 highlighted how designing anywhere — including outdoors — opened new possibilities, such as creating projects in natural settings. The immersive experience was widely appreciated, with features like depth occlusion enhancing realism. As P10 noted, “The occlusion made everything feel so much more immersive, like the objects were truly part of the room.” The engaging and enjoyable nature of the system was also mentioned, with P3 stating, “Using the system was genuinely fun and engaging.”
Intuitive and Natural Controls
The hand-based interactions and intuitive controls stood out as significant strengths. P6 described them as “human-nature,” explaining, “Locking axes was so simple, just a left-hand finger pinch and moving along the axis I wanted.” Similarly, the snap-to-grid feature was highly valued for its precision, with P2 commenting, “The snap-to-grid feature was a game-changer for precise alignments.” These intuitive mechanisms allowed participants to focus on creativity without being hindered by complex operations.
Physical Demands and Ergonomics
While the controls were intuitive, several participants expressed concerns about the physical demands of prolonged use. P1 mentioned, “The physical demand was rather tiring,” and P2 noted, “It was a little tiring to have my arm raised for every action.” These observations suggest that extended use could lead to fatigue, especially when designing complex models.
Accessibility and Future Adoption
TinkerXR was widely commended for its accessibility, even among novice users. P6 shared, “As someone completely new to AR, 3D modeling, and 3D printing, I was surprised by how easy it was to learn.” This ease of use motivated participants to envision future applications, such as designing custom objects for their homes or gardens. P4 expressed excitement about using the system for personal projects, stating, “Even in my garden, I am already thinking about designing a birdhouse and some tree decorations.” Most participants indicated they would continue using the tool and recommend it to others, with P8 emphasizing its convenience: “Starting the print directly from the headset saves materials because I don’t have to redo prints.”
Feature Expansion and Precision Tools
Participants identified areas where additional features could enhance the workflow. Several requested better tools for manipulating small objects, with P7 suggesting a “magnifying glass tool for designing small objects.” Building on the AR ruler feature, on-screen measurements were another commonly desired functionality, as P1 noted, “Length values of the edges appearing next to the objects, just like in Tinkercad, would help.” Other recommendations included importing external 3D models and freely adjusting grid sizes. P4 also proposed more direct interaction methods, such as “grabbing objects directly with our hands.”
Overall, the feedback highlights TinkerXR’s strengths in providing an immersive, intuitive, and accessible design experience, especially for novices. However, challenges like solving the problem of physical fatigue on extended usage should be solved in the future. Addressing functionalities like small object manipulation and hand-grabbing the objects would also further enhance its usability.
7. Discussion
In this section, we evaluate the effectiveness and future potential of TinkerXR. We discuss its suitability for both novice and expert users and express limitations related to handling complex 3D geometries. We also address the impact of the AR form factor on user experience and the limitations of hand gesture recognition. Additionally, we anticipate future advancements that could improve the system’s capabilities and user interaction.
Novice vs. Expert Usage
Our tool was designed with novices in mind, providing an intuitive interface that simplifies both the design and fabrication of models. However, we envision that TinkerXR can also be beneficial to experts. Practitioners and professional designers can use it for rapid prototyping, concept development, and interactive presentations. The ease of use and flexibility of our system can streamline workflows, allowing experts to focus on more complex design tasks without being bogged down by technical logistics. In the future, we believe that our system can be adopted by educational institutions for teaching design principles, by industries for quick mock-ups or collaborative sessions, and by hobbyists seeking to explore 3D modeling and fabrication.
3D Geometry Complexity
In Section 5, we demonstrated our tool’s usefulness through various design examples, each showing a different use case. These objects featured distinct geometries, ranging from small details to larger-scale structures. For instance, the candy bowl model in Figure 14d is made of 32 curved lines, each made of 8 cuboids, adding up to a total of 256 basic building blocks (if none of the blocks are combined) and 2048 vertices. Our system supports up to 300 basic building blocks and a total number of vertices of 20,000 in real time. However, we observed a noticeable lag when exceeding these numbers, especially in combination operations. We anticipate this limitation can be addressed with future headsets offering higher processing power. Alternatively, more powerful headsets, e.g., Varjo XR-4141414https://varjo.com/products/xr-4/, Accessed: 4.08.2024, which is connected to a desktop computer, can support compute-intensive tasks.
While the system can create a variety of complex geometries, certain tasks remain challenging due to the lack of more advanced tools such as extrusion of 3D geometries from 2D blueprints (e.g., sketched shapes, text), parametric curves, lofting, filleting, freeform design tools similar to those in Autodesk 3ds Max [^15] https://www.autodesk.com/products/3ds-max, Accessed: 4.08.2024, or a collaborative structure allowing multiple users to work on the same project similar to Gravity Sketch or adding physical engagement with local users and their surroundings similar to HoloBots (Ihara et al., 2023). These features were not implemented, as our system targets novices and prioritizes simplicity. While it is arguable that adding such features inherently increases complexity, our design choices focus on maintaining an intuitive user experience.
We only implemented the features that were required for realizing the artifacts illustrated in Section 5 and objects with similar structural complexity, which were inspired by Tinkercad repositories and related work. In the future, because we open-source our work, we believe that the community can add more features to enhance its capabilities. By fostering a collaborative development environment, we hope to see our system evolve to meet the needs of users of every proficiency, ultimately broadening its applicability and usability.
Scaling to Other Fabrication Methods
We believe that TinkerXR has the potential to adapt to other digital fabrication methods such as molding (Serrano et al., 2022; Valkeneers et al., 2019), CNC machining (Tran O’Leary et al., 2023; Tran O’Leary and Peek, 2019), or laser cutting (Dogan et al., 2023a, 2021) in the future. However, we acknowledge that this adaptation goes beyond simple file format conversion. Each method introduces unique design constraints, such as draft angles for molds, toolpath limitations for CNC milling, and material thickness or 2D geometry constraints for laser cutting. These factors would require the system to integrate CAM tools to guide users in adhering to method-specific constraints.
In this context, Roumen et al. (Baudisch et al., 2019) provide inspiration for adapting 3D design for laser cutting, although it is not a general-purpose solution. Exploring ways to integrate such techniques into TinkerXR could enable users to understand fabrication-specific constraints in real time, which we consider an exciting avenue for future work.
AR Form Factor
MR headsets can be bulky; however, in our user study, participants did not report major problems due to the form factor. We used the Quest 3 because it is lightweight, making it easy to use. In the future, we imagine AR technology becoming more ubiquitous while offering a more streamlined experience with the advent of AR glasses, such as Meta Orion [^16] https://about.meta.com/realitylabs/orion/, Accessed: 20.01.2025. Future devices are likely to be more compact, resembling ordinary eyewear, thus enhancing user comfort and accessibility (Dogan, 2024). These advancements could lead to broader adoption and more immersive systems.
Hand Gestures
While we are constrained by Meta’s hand gesture APIs, which are known to improve over time [^17] https://www.techradar.com/computing/virtual-reality-augmented-reality/your-meta-quest-3-is-getting-a-hand-tracking-upgrade-that-could-unlock-foot-tracking, Accessed: 5.08.2024, there are still issues such as improper functionality in dimly lit areas or limited precision of tracking compared to controllers or the system being exclusive to Meta headsets.
8. Conclusion
In this paper, we presented TinkerXR, a new 3D modeling and fabrication tool through AR, designed for novices. Our approach lets users design artifacts with a simple primitive geometry manipulation CSG design logic, with intuitive hand gesture control schemes, on the intended space of the artifact, and with seamless integration with the 3D printing process. We showed the various features present in TinkerXR from selecting the workspace to design and fabrication. Using these features, we presented a wide variety of creations for different scenarios, each varying in complexity and practical applications. A user study revealed that the system was effective for novices, and more immersive than Tinkercad, while offering a comparable level of functionality and simplicity. We hope that our project and its open-source codebase will stimulate further exploration within communities engaged in HCI, design, and fabrication.
Acknowledgements.
This work was supported by the INVERSE Project under Grant No. 101136067, funded by the European Union.